A ball (m=0.8 kg) falls to the floor from a height of 2 m and bounces back to a height of 1.2 m. Calculate the amount of energy lost due to air drag and collision with the ground.


Could you please explain it in detail?


1
Expert's answer
2021-06-14T08:51:39-0400

First, find the initial energy, the initial potential energy:


"E_i=mgh_i=0.8·9.8·2=15.68\text{ J}."

The final energy (after rebounding):


"E_f=mgh_f=0.8·9.8·1.2=9.41\text{ J}."

How much was lost? Right:


"E_L=E_i-E_f=15.68-9.41=6.27\text{ J}."